# WFR-FM: Simulation-Free Dynamic Unbalanced Optimal Transport

### v1

@@ Abstract
The Wasserstein–Fisher–Rao (WFR) metric extends dynamic optimal transport (OT) by coupling displacement with change of mass, providing a principled geometry for modeling unbalanced snapshot dynamics. Existing WFR solvers, however, are often unstable, computationally expensive, and difficult to scale. Here we introduce WFR Flow Matching (WFR-FM), a simulation-free training algorithm that unifies flow matching with dynamic unbalanced OT. Unlike classical flow matching which regresses only a transport vector field, WFR-FM simultaneously regresses a vector field for displacement and a scalar growth rate function for birth–death dynamics, yielding continuous flows under the WFR geometry. Theoretically, we show that minimizing the WFR-FM loss exactly recovers WFR geodesics. Empirically, WFR-FM yields more accurate and robust trajectory inference in single-cell biology, reconstructing consistent dynamics with proliferation and apoptosis, estimating time-varying growth fields, and applying to generative dynamics under imbalanced data. It outperforms state-of-the-art baselines in efficiency, stability, and reconstruction accuracy. Overall, WFR-FM establishes a unified and efficient paradigm for learning dynamical systems from unbalanced snapshots, where not only states but also mass evolve over time.
@@ 2 Related Works
Single-cell Trajectory Inference. Many algorithms have been proposed for single-cell trajectory inference in the multiple-time-points setup, including static optimal transport maps Schiebinger et al. (2019); Klein et al. (2025); Halmos et al. (2025); Banerjee et al. (2024), continuous formulations via neural ODEs Tong et al. (2020); Huguet et al. (2022); Zhang et al. (2024a); Sha et al. (2024); Tong et al. (2023); Gu et al. (2025); Choi and Choi (2024), and stochastic variants Neklyudov et al. (2023; 2024); Albergo et al. (2023); Zhu et al. (2024); Maddu et al. (2024); Yeo et al. (2021); Chizat et al. (2022); Ventre et al. (2023); Lavenant et al. (2024); Shi et al. (2024); Koshizuka and Sato (2023); Bunne et al. (2023); Chen et al. (2022b); Jiang and Wan (2024); Zhang et al. (2025a); Sun et al. (2025). More recently, flow matching has been applied in this domain to improve scalability and training stability (Tong et al., 2024a; Rohbeck et al., 2025; Klein et al., 2024; Tong et al., 2024b; Lee et al., 2025; Kapusniak et al., 2024; Atanackovic et al., 2025). The biological importance of proliferation and apoptosis has further motivated extensions to unbalanced settings (Neklyudov et al., 2023; 2024; Eyring et al., 2024; Wang et al., 2025). Notably, VGFM (Wang et al., 2025) explicitly incorporating growth estimation into the flow-matching framework. However, existing unbalanced approaches either rely on static approximations or require costly ODE simulations, or they lack a truly simulation-free flow matching framework that can jointly recover velocity and growth in single-cell dynamics.
@@ 3 Mathematical Background
When conditioning pairs $(\bm{x}_{0},\bm{x}_{1})$ are drawn from an OT coupling $\gamma(\bm{x}_{0},\bm{x}_{1})$, the resulting objective recovers the dynamic OT flow. Thus, the design of conditional path, conditional loss, and coupling (i.e. conditions) jointly determines whether the learned dynamics match classical OT geodesics or more general bridges.
@@
The relation between $(\gamma_{0},\gamma_{1})$ and $\gamma$ has been studied previously under general condition (Liero et al., 2018). In this work, we also give a straightforward proof to the case of the WFR problem that the optimal semi-coupling can be constructed from the OET coupling $\gamma$ as following.
@@ 4.3 Coupling in WFR-FM solves the dynamical WFR problem
where $\bm{\omega}_{0}$ and $m_{t}(\bm{x}_{0},\bm{x}_{1})$ are defined as (3.5) and (3.6) with boundary condition $m_{0}(\bm{x}_{0},\bm{x}_{1})=1,\ m_{1}(\bm{x}_{0},\bm{x}_{1})=\frac{\gamma_{1}(\bm{x}_{0},\bm{x}_{1})}{\gamma_{0}(\bm{x}_{0},\bm{x}_{1})}$. We set $m_{0}=1$ because $m_{t}$ describes the mass evolution relative to the initial mass along each trajectory, and fixing the initial mass to 1 provides a consistent reference. Under this setting, we can show that the marginal boundary measures approach $\mu_{0}$ and $\mu_{1}$ respectively as $\sigma\to 0$, and the induced measure path follows the WFR geodesic.
@@
The proof is left to Appendix A.5. Note that even in balanced cases where $m_{0}=m_{1}=m$, the mass trajectory induced by travelling Dirac (3.5) under WFR metric is still not constant, but instead being a quadratic function of time $m(t)=2m(1-\frac{1}{\sqrt{1+\tau^{2}}})t(1-t)+m$. To keep the mass conserved during the whole process, one needs to let the penalization of growth $\delta\to\infty$ (3.3) to ensure no growth. This limiting case of WFR-FM, which is consistent with balanced dynamical OT, is characterized by the following proposition, whose proof is left to Appendix A.6.
@@ 5 WFR-FM Workflow for Trajectory Inference
Multi-time point WFR Formulation. Given samples from distributions $\mu_{t_{i}}$ at $K+1$ discrete time points, $t=t_{0},t_{1},\ldots,t_{K}$, our goal is to reconstruct the continuous evolution of these distributions. To this end, we solve the WFR problem, which entails minimizing the WFR action functional subject to the constraint that $\rho(\bm{x},t)$ coincides with the observed distributions $\mu_{t_{i}}$ at the specified time points. Formally,
@@
Equivalently, the multi-time problem reduces to solving WFR between successive time points, guarateed by the following claim proved in Appendix A.7.
@@
In practical training, we impose that the time-dependent vector fields and growth rates are shared across different successive time intervals to improve the generalization capability.
@@ 6 Experiment Results
In addition, we conducted a sensitivity analysis of key hyperparameters, including the penalty coefficient for the growth term $\delta$ (Table 5) and the batch size in the mini-batch WFR-OET (Table 8), and found that WFR-FM is robust to these choices.
Evaluation of Learned Growth Dynamics (Q5) To explicitly validate whether WFR-FM captures the underlying birth-death dynamics rather than merely matching marginal distributions, we evaluate the accuracy of the learned growth rate function $g_{\phi}(x,t)$. Since the ground truth growth rate is typically unavailable in real-world biological data, we conduct this evaluation on the Simulation Gene dataset, where the dynamics are governed by known SDEs. We calculate the Pearson correlation coefficient ($g_{\text{corr}}$) between the predicted growth rate and the ground truth growth rate $g=\alpha_{g}\frac{X_{2}^{2}}{1+X_{2}^{2}}$ on the data points from the dataset. As shown in Table 4, WFR-FM achieves the highest correlation, outperforming baseline methods. This demonstrates that WFR-FM provides a reliable estimation of the cell proliferation and apoptosis rates.
Table 4: Correlation between the predicted and ground truth growth rates ($g_{\text{corr}}$) on the Simulation Gene dataset. Best results are in bold.
Method Correlation ($g_{\text{corr}}$)
TIGON 0.9705
DeepRUOT 0.9688
Var-RUOT 0.9214
Action Matching 0.5851
WFR-FM (Ours) 0.9913
@@ 7 Conclusion, Limitation and Discussion
In this work, we introduced WFR-FM, a framework that unifies the WFR metric with flow matching to address dynamic unbalanced OT in a simulation-free manner. By jointly learning velocity fields and growth rates, WFR-FM provides the principled formulation of unbalanced flow matching with theoretical guarantees, exactly solving the WFR problem. Experiments on scRNA-seq datasets demonstrate its effectiveness in reconstructing single-cell dynamics with proliferation and apoptosis, highlighting its potential as a robust tool for trajectory inference in biology.
A limitation shared with other methods combining OT and FM is that WFR-FM relies on solving a static OT problem, which can be costly for very large datasets. To address this, we adopt approximation strategies such as mini-batch OT and entropy-regularized Sinkhorn iterations. Incorporating uncertainty quantification in WFR-FM could also enhance the performance on noisy single-cell datasets. While our focus has been on WFR geometry, the framework established is indeed general: as long as the static OT problem can be efficiently solved and the Dirac-to-Dirac path admits a closed-form solution (e.g., when linear growth penalties are chosen, as in Chizat et al. (2018b)), our approach can be extended to other unbalanced transport functionals. We anticipate that this generality will enable new classes of flow-matching methods, broadening their applications across scientific domains where mass is not conserved or other machine learning application scenarios with unbalanced datasets.
@@
We have taken several measures to ensure the reproducibility of our work. All theoretical results are accompanied by complete proofs provided in the Appendix. Detailed descriptions of data processing steps are also included in the Appendix to facilitate replication of our experiments. The source code will be made publicly available upon acceptance.

### v2

@@ Abstract
The Wasserstein–Fisher–Rao (WFR) metric extends dynamic optimal transport (OT) by coupling displacement with change of mass, providing a principled geometry for modeling unbalanced snapshot dynamics. Existing WFR solvers, however, are often unstable, computationally expensive, and difficult to scale. Here we introduce WFR Flow Matching (WFR-FM), a simulation-free training algorithm that unifies flow matching with dynamic unbalanced OT. Unlike classical flow matching which regresses only a transport vector field, WFR-FM simultaneously regresses a vector field for displacement and a scalar growth rate function for birth–death dynamics, yielding continuous flows under the WFR geometry. Theoretically, we show that minimizing the WFR-FM loss exactly recovers WFR geodesics. Empirically, WFR-FM yields more accurate and robust trajectory inference in single-cell biology, reconstructing consistent dynamics with proliferation and apoptosis, estimating time-varying growth fields, and applying to generative dynamics under imbalanced data. It outperforms state-of-the-art baselines in efficiency, stability, and reconstruction accuracy. Overall, WFR-FM establishes a unified and efficient paradigm for learning dynamical systems from unbalanced snapshots, where not only states but also mass evolve over time. The Python code is available at https://github.com/QiangweiPeng/WFR-FM.
@@ 2 Related Works
Single-cell Trajectory Inference. Many algorithms have been proposed for single-cell trajectory inference in the multiple-time-points setup, including static optimal transport maps Schiebinger et al. (2019); Klein et al. (2025); Halmos et al. (2025); Banerjee et al. (2024), continuous formulations via neural ODEs Tong et al. (2020); Huguet et al. (2022); Zhang et al. (2024a); Sha et al. (2024); Tong et al. (2023); Gu et al. (2025); Choi and Choi (2024), and stochastic variants Neklyudov et al. (2023; 2024); Albergo et al. (2023); Maddu et al. (2024); Yeo et al. (2021); Chizat et al. (2022); Ventre et al. (2023); Lavenant et al. (2024); Shi et al. (2024); Koshizuka and Sato (2023); Bunne et al. (2023); Chen et al. (2022b); Jiang and Wan (2024); Zhang et al. (2025a); Sun et al. (2025). More recently, flow matching has been applied in this domain to improve scalability and training stability (Tong et al., 2024a; Rohbeck et al., 2025; Klein et al., 2024; Tong et al., 2024b; Lee et al., 2025; Kapusniak et al., 2024; Atanackovic et al., 2025). The biological importance of proliferation and apoptosis has further motivated extensions to unbalanced settings (Neklyudov et al., 2023; 2024; Eyring et al., 2024; Wang et al., 2025). Notably, VGFM (Wang et al., 2025) explicitly incorporating growth estimation into the flow-matching framework. However, existing unbalanced approaches either rely on static approximations or require costly ODE simulations, or they lack a truly simulation-free flow matching framework that can jointly recover velocity and growth in single-cell dynamics.
@@ 3 Mathematical Background
When conditioning pairs $(\bm{x}_{0},\bm{x}_{1})$ are drawn from an OT coupling $\gamma(\bm{x}_{0},\bm{x}_{1})$, the resulting objective recovers the dynamic OT flow. Thus, the conditional path, conditional loss, and coupling (i.e. conditions) designs jointly determine if learned dynamics match OT geodesics or general bridges.
@@
The relation between $(\gamma_{0},\gamma_{1})$ and $\gamma$ has been studied previously under general condition (Liero et al., 2018). In this work, we also give a straightforward proof (see Appendix A.1) for the WFR problem that the optimal semi-coupling can be constructed from the OET coupling $\gamma$ as following.
@@ 4.3 Coupling in WFR-FM solves the dynamical WFR problem
where $\bm{\omega}_{0}$ and $m_{t}(\bm{x}_{0},\bm{x}_{1})$ are defined as (3.5) and (3.6) with boundary condition $m_{0}(\bm{x}_{0},\bm{x}_{1})=1,\ m_{1}(\bm{x}_{0},\bm{x}_{1})=\frac{\gamma_{1}(\bm{x}_{0},\bm{x}_{1})}{\gamma_{0}(\bm{x}_{0},\bm{x}_{1})}$. We set $m_{0}=1$ because $m_{t}$ describes the mass evolution relative to the initial mass along each trajectory. Under this setting, we can show that the marginal boundary measures approach $\mu_{0}$ and $\mu_{1}$ respectively as $\sigma\to 0$, and the induced measure path follows the WFR geodesic (proof in Appendix A.5).
@@
Notably, even for balanced masses $m_{0}=m_{1}=m$, the WFR mass trajectory $m(t)=2m(1-\frac{1}{\sqrt{1+\tau^{2}}})t(1-t)+m$ remains non-constant. Mass conservation is only recovered as the growth penalty $\delta\to\infty$, leading to the balanced OT limit characterized below (proof in Appendix A.6).
@@ 5 WFR-FM Workflow for Trajectory Inference
Multi-time point WFR Formulation. Given samples from distributions $\mu_{t_{i}}$ at $K+1$ discrete time points, $t=t_{0},t_{1},\ldots,t_{K}$, our goal is to reconstruct the continuous evolution of these distributions. To this end, we solve the WFR problem, which entails minimizing the WFR action functional subject to the constraint that $\rho(\bm{x},t)$ coincides with the observed distributions $\mu_{t_{i}}$ at the specified time points.
@@
The multi-time problem (5.1) reduces to a sequence of WFR steps, as established by the following proposition (proof in Appendix A.7). In practice, we share vector fields and growth rates across all time intervals to enhance generalization.
@@ 6 Experiment Results
Sensitivity analysi confirms WFR-FM’s robustness to key hyperparameters, specifically the growth penalty $\delta$ (Table 5) and batch size for mini-batch WFR-OET (Table 8).
Evaluation of Learned Growth Dynamics (Q5) To confirm WFR-FM captures birth-death dynamics beyond marginal matching, we evaluate the learned growth rate $g_{\phi}(x,t)$ on the Simulation Gene dataset, where ground truth is provided by known SDEs. We calculate the Pearson correlation coefficient ($g_{\text{corr}}$) between the predicted growth rate and the ground truth growth rate $g=\alpha_{g}\frac{X_{2}^{2}}{1+X_{2}^{2}}$ on the data points from the dataset. Table 4 shows that WFR-FM outperforms baselines, validating its ability to reliably estimate growth rates.
Table 4: Growth rate correlation ($g_{\text{corr}}$) on Simulation Gene dataset. Best results are in bold.
Method TIGON DeepRUOT Var-RUOT Action Matching WFR-FM (Ours)
Correlation ($g_{\text{corr}}$) 0.9705 0.9688 0.9214 0.5851 0.9913
@@ 7 Conclusion, Limitation and Discussion
In this work, we introduce WFR-FM, a framework unifying the WFR metric with flow matching for simulation-free, dynamic unbalanced OT. By jointly learning learning velocity fields and growth rates, WFR-FM provides a principled formulation with theoretical guarantees that exactly solves the WFR problem. Extensive experiments demonstrate its effectiveness in reconstructing dynamics with proliferation and apoptosis, establishing it as a robust tool for biological trajectory inference.
Like other OT-FM methods, WFR-FM relies on solving a static OT problem, which can be costly for very large datasets. To address this, we adopt approximation strategies such as mini-batch OT and entropy-regularized Sinkhorn iterations. Incorporating uncertainty quantification in WFR-FM could also enhance the performance on noisy single-cell datasets. While our focus has been on WFR geometry, the framework established is indeed general: as long as the static OT problem can be efficiently solved and the Dirac-to-Dirac path admits a closed-form solution (e.g., when linear growth penalties are chosen, as in Chizat et al. (2018b)), our approach can be extended to other unbalanced transport functionals. We anticipate that this generality will enable new classes of flow-matching methods, broadening their applications across scientific domains where mass is not conserved or other machine learning application scenarios with unbalanced datasets.
@@
We have taken several measures to ensure the reproducibility of our work. All theoretical results are accompanied by complete proofs provided in the Appendix. Detailed descriptions of data processing steps are also included in the Appendix to facilitate replication of our experiments. The source code is available at https://github.com/QiangweiPeng/WFR-FM.